OpenStudy (anonymous):

Using complete sentences, explain how to graph y = one thirdx + 4 using the slope-intercept method.

OpenStudy (anonymous):

y = (1/3) x + 4 Start at x=0, y=4, and draw your line through that point and a point at x = 3, y = (1/3)(3) + 4 = 5. You can continue this line in either direction as far as you need to.

OpenStudy (anonymous):

The slope is 1/3, 1 up for y for every 3 over for x. The intercept is b=4, which is why we started with x=0, y=4.

OpenStudy (texaschic101):

in y = mx + b form, the m position is the slope and the b position is the y intercept. So you have a slope of 1/3 and a y intercept of (0,4) to find the x intercept, sub in 0 for y y = 1/3x + 4 0 = 1/3x + 4 -1/3x = 4 x = 4/(-1/3) x = 4 * -3 x = -12 x intercept is (-12,0) plot your points (0,4) and (-12,0). Connect them with a solid line.